# Johannes Matzat        Aiko Schmeißer

### v4

@@ 6.2 Ideological versus Strategic Motives
As discussed in Section 2.2, the literature considers campaign contributions as driven by either expressive (ideological) or investment-oriented (strategic) motives. While a large strand of work on money in politics finds that individual donors allocate their contributions primarily based on ideological preferences (Stone and Simas, 2010; Bonica, 2014, 2018; Barber, 2016a, b), more recent research has found that parts of employees’ donations are also related to their employers’ strategic interests (Stuckatz, 2022; Teso, forthcoming). Along that line, the observed shift in managers’ donations from Republican to Democratic candidates could reflect, for example, a strategic signaling of goodwill and understanding of unions’ and workers’ interests instead of a true shift in managers’ preferences. Moreover, unionization could change workers’ efforts to support candidates in exchange for future legislative votes on worker-friendly policies. While it is difficult to fully disentangle donors’ motives, in the following we present several pieces of evidence that all point toward changes in ideological preferences being the main driver of our results.
@@
Differentiating candidates by within-party ideology. To examine the ideological patterns in campaign contributions in more detail, we study the changing support for ideologically different candidates within the same party. Kuziemko et al. (2023) document a growing influence of conservative factions in the Democratic party that favor more educated voters and focus less on predistribution economic policies, such as higher minimum wages, protectionism and strong labor unions. To analyze how unionization affects the support of different party factions, we make use of Bonica’s (2014) CF scores that assign each recipient an ideal point along a liberal-conservative scale. Democratic candidates are categorized as “moderate” versus “liberal” if their CF score lies above the median CF of all Democrats observed in our sample of matched contributions. Similarly, we distinguish between “moderate” and “conservative” Republicans using the median Republican CF score. Table 5 shows the DiD results for the effect on the amount contributed to each of the candidate types. Considering first all employees jointly, we see strong differences in the effects of unionization by the within-party ideological positions of candidates. Unionization significantly increases employees’ support for the most liberal Democrats and decreases support for the most conservative Republicans. In contrast, contributions to moderate Democrats or Republicans are not significantly affected. These results are similar when we focus on donations from managers only, and also for workers the increased support for Democrats is more pronounced for more liberal Democrats. Overall, our results do not seem to be driven by a mere signal of affiliating with the Democratic party but instead reflect a shift in contributions between clearly distinguishable conservative and liberal candidates.
@@
Differentiating candidates by union support. We continue by examining whether the partisan shift in employee contributions can be solely explained by a shift from candidates not supported to candidates supported by the union organization. To identify union-supported candidates, we have matched donations from PACs associated with the union organization that was on the ballot in a given union election (including PACs of local union branches). Appendix Table A.11 reports for each union organization in our sample the share of contributions to Democratic (as opposed to Republican) candidates. On average, union PACs give 94% of their donations to Democrats, which demonstrates the strong alliance between labor unions and the Democratic Party (Dark, 2001). In Table 6, we use this information to estimate the effects of unionization on employees’ contribution amounts to candidates supported and not supported by the respective union in a cycle. Unions may encourage their members to contribute to the same candidates as their PAC, either directly through providing information about these candidates or indirectly through employees contributing to the union PAC and observing which candidates are supported by the union PAC. While the estimated coefficients for the effect of unionization on employees’ contribution amounts to union-supported candidates are positive, they are small and insignificant. [^40] When taking the difference in the amounts going to union-supported and non-union-supported candidates, the effect of unionization is significant only for contributions from managers and only at the 10% level (not shown). Moreover, we find a shift in contributions from Republican to Democratic candidates that remains significant even after conditioning on whether or not candidates are supported by the union. In other words, among both union-supported and non-union-supported candidates we observe significant effects on contribution amounts to Democratic relative to Republican candidates. These results suggest a limited role for unions’ direct influence in channeling employees’ contributions toward candidates in which the union has strategic interests. Instead, unionization appears to cause a more general leftward shift in employees’ donation behavior.
@@
Contributions to PACs. In Table 2, we have shown that contributions to PACs account for a large share of political contributions from employees. Besides changing employees’ support of specific candidates, unionization may also affect the donations to intermediary committees with varying ideological positions. Most evidently, unions may solicit contributions from their members to their own PACs. Table 7 reports DiD estimates for the effect of unionization on employees’ donations to party PACs and interest group PACs, where the latter are further disaggregated into labor union, membership organization, corporate, and trade association PACs. Besides considering the total amount given to these committees, we also measure partisan support by the difference in contribution amounts to Democratic versus Republican PACs. For interest group PACs, party affiliation is determined from the recipients of the PAC’s own campaign contributions. [^41] To track contributions that PACs donate themselves, we exploit that Bonica (2019) has matched recipient identifiers to contributor identifiers for recipients’ own contributions. Based on the matched outgoing contributions from PACs, we define an interest group PAC as “Democratic” (“Republican”) if more (less) than 50% of its campaign contributions goes to Democratic candidates in a given election cycle. Considering first the contributions from all employees of an establishment to party PACs, the results mimic those for candidate contributions. While there is no effect on total amounts, unionization leads to a significant shift from Republican to Democratic committees. Among interest group PACs, there is a significantly positive effect on donations to union PACs and a significantly negative effect on donations to corporate PACs. When distinguishing between donations from workers and managers, results differ somewhat. For workers, we see a significant increase in the total amounts donated to both party and interest group PACs, which implies that unions are successful in mobilizing PAC contributions from workers. The increase in donations appears to be driven by unions and membership organizations, pointing toward increased support for labor and civil society interest groups. In contrast to our results on candidates, however, we do not see a significant shift across party affiliations. For managers, the results are similar to those on candidate contributions. While there is no effect on overall party PAC spending, managers increasingly donate to Democratic rather than Republican PACs. Moreover, donations to corporate PACs drop, which highlights that unionization can decrease managers’ support for business interest groups. Overall, these results match with the observed pro-liberal shift in workers and managers’ contributions to political candidates.
@@
Contributions from firm PACs. Finally, we directly examine whether unionization changes strategic donations of unionized firms. A growing body of evidence shows that corporations direct contributions from their PACs in ways that are consistent with an attempt to buy access to relevant politicians (Kalla and Broockman, 2016; Powell and Grimmer, 2016; Fouirnaies and Hall, 2018). For each union election in our estimation sample, we link candidate contributions from their firm’s PAC and aggregate them at the firm level. [^42] We link firm PAC contributions using a fuzzy match of the firm name. Given that the location of the firm PAC may differ from that of the establishment, we do not match on the commuting zone of the establishment. For 1,902 out of the 6,603 establishments in our estimation sample, we observe at least one firm PAC contribution over our observation window. In Appendix Table A.12, we report results for the full sample of establishments when considering unmatched observations as zeros, as well as for the reduced sample of establishments with at least one matched firm PAC contribution. The results of our DiD model, reported in Appendix Table A.12, show that unionization does not significantly affect firm PACs’ total contribution amounts or their partisan composition. We conclude that the leftward shift in employees’ donations does not seem to coincide with a corresponding shift in employers’ strategic donations.
6.3 Union Engagement
Unions are not a uniform political force; instead, they are heterogeneous and evolving organizations that may vary in their ability to politically persuade employees (Ahlquist and Levi, 2013; Kim and Margalit, 2017). In the following, we examine how the effects of unionization on employees’ campaign contributions vary with unions’ engagement, in terms of both their political activities and their ability to secure a collective bargaining agreement.
Right-to-Work laws. An important institutional factor influencing the political power of unions is state-level Right-to-Work legislation. RTW laws make paying union dues optional. This means that even if workers in a unionized workplace benefit from union representation and the negotiated bargaining agreements, they do not have to become union members and pay fees. Feigenbaum et al. (2018) study in detail the political consequences of RTW laws. They provide evidence that RTW legislation pressures union revenues and forces unions to reallocate scarce resources away from political activities (such as lobbying, voter mobilization, candidate recruitment, or campaign contributions) toward membership recruitment activities. In terms of electoral consequences, the passage of RTW laws is found to reduce voter turnout and vote shares for Democratic candidates at the regional level.
@@ 6.3 Union Engagement
We complement this analysis by studying how RTW laws moderate the effect of unionization on employees’ campaign contributions at the establishment level. It is the unionized workplace where unions are directly connected to employees and where RTW laws may thus have a large impact on unions’ political influence. To test this hypothesis, we split our estimation sample based on whether or not the union election takes place in a state that has an RTW law in force at the time of the election. Results are presented in Table 8, Panels A.1 and A.2. In states without RTW laws, we see significantly positive effects of unionization on support for Democratic over Republican candidates, while for RTW states the coefficients are smaller and not significant. This is true for all employees, as well as for workers and managers separately. Thus, reduced political activity of unions under RTW legislation seems to weaken their ability to channel campaign contributions from employees at unionized workplaces.
Collective bargaining contracts. We also examine the role of union strength at the bargaining table, measured by whether the union successfully negotiates a collective bargaining agreement with the employer after winning the union election. For this, we compile and merge data on contract negotiation notices from the Federal Mediation and Conciliation Service (FMCS). [^43] We use contract data from Holmes (2006) for 1985 - 2003 (link) and from Forest Gregg for 2004 - 2020 (link). The data includes notices of contract expirations that are provided to the FMCS for it to be able to prepare mediation services for the renegotiation of a contract. We consider contract expiration notices in the 5 years following the union election. In our sample, only 48% of the won elections report a bargaining agreement. In Panels B.1 and B.2 of Table 8, we split our treatment group in won union elections with and without a reported bargaining contract. [^44] In both cases, we compare the won union elections to the control group of all lost union elections. For union certifications that result in a collective bargaining agreement, we find strong positive effects on campaign donations to Democratic over Republican candidates, while there is no significant effect on won elections without a bargaining agreement.
Importantly, we find this heterogeneity for contributions from both workers and managers. For workers, the results suggest that organizing alone is not sufficient to give rise to a political impact. Instead, it appears that the union must also deliver on the bread and butter issues in order to be able to politically influence their members. Managers, who are not part of the bargaining unit, do not receive direct economic benefits from a collective bargaining agreement. Still, we also observe a larger shift in managers’ donations from Republican to Democratic candidates in unionized establishments where they agree with the union on a collective bargaining agreement. This result is consistent with the idea that positive, continuous contact and perspective taking can explain why managers align their political positions with workers and the union.
6.4 Quantitative Interpretation
This paper leverages campaign contribution patterns as a proxy for ideological positions of employees. We interpret a shift in donations from Democratic to Republican candidates as informative independent of the overall amount of money donated by a given individual. Sections 2.2 and 6.2 have discussed the plausibility of interpreting shifts in campaign contributions as shifts in ideological positions.
In this section, we offer an alternative perspective on our results that focuses solely on the amount of money shifted by unionization. Campaign contributions are essential for winning elections in the U.S, funding activities such as political advertising, fundraising, and campaign staff (Open Secrets, 2023). The average amount spent by winners of House elections was \$777,000 in 1986, which has increased to \$1.47 million by 2014 (Persily et al., 2018). Although acess to considerable campaign funding is not seen as sufficient, it is deemed necessary to win elections (Jacobson, 2015). This section quantifies the total shifts in campaign contributions likely caused by unions, by extrapolating our findings to all unionized establishments.
We start by estimating the amount contributed by employees at unionized establishments. We first calculate the cycle-specific average amount donated per voter in unionizing establishments in our sample and then multiply it by the number of workers represented by unions during that cycle. Summing up over all cycles from 1986 to 2010, employees at unionized establishments have donated \$4.31 billion. [^45] Formally, $y=\sum_{c}\left(\frac{\textrm{Amount}_{c}}{\textrm{Voters}_{c}}\times\textrm{\#Represented}_{c}\right)$, where $c$ denotes cycles. We sum up individuals’ donations to both candidates and PACs. The magnitude is sizeable and highlights that individual donations complement donations through union PACs, which amount to \$5.88 billion for the unions in our sample [^46] We have excluded contributions to union PACs from the total individual donations to avoid double counting.. In total, \$10.21 billion are under direct or indirect influence of unions.
To quantify the magnitude of the influence of unions on the partisan composition of campaign contributions, we repeat the exercise for our main causal estimate. We first translate the percentage change to dollar amounts using mean contribution amounts per establishment. This results in an average shift of \$250.86 from Republican to Democratic recipients (candidates or PACs) per establishment and cycle. In a second step, we derive the cycle-specific average shift per voter and multiply it by the number of represented employees. Aggregating over all cycles, we obtain the total shift in donations caused by unions: \$511.3 million are shifted from Republican toward Democratic candidates. [^47] We re-estimate our main specification after aggregating candidate and PAC contributions, using the difference of IHS-transformed contribution amounts to Democratic minus Republican recipients as the outcome. The obtained point estimate is 0.270 (SE: 0.086). Multiplying with the mean contribution amount to Republican candidates per establishment gives 0.270 $\times$ \$929.11 = \$250.86, which represents the average shift in donations from Republican to Democratic recipients. We extrapolate this sum to all unionizing establishments as follows: ${y=\sum_{c}\left(\frac{250.86}{\textrm{Voters}_{c}}\times\textrm{\#Represented}_{c}\right)}$. Contrasting the \$511.3 million to the cost of running successfully for a seat in the House of Representatives (which lies between \$777,000 and \$1.47 million (Persily et al., 2018)), highlights the sizeable influence of unions via their impact on donations of employees at unionized establishments.
Our results contribute to a better understanding of the political consequences brought about by the structural decline in unionization. The share of workers represented by unions has continuously declined from 23.3% in 1983 to 13.1% in 2010 (BLS, 2023). To quantify the political implications of the disappearance of unions over our sample period, we contrast our results with estimates based on a counterfactual scenario in which the rate of workers represented by unions would have remained at its 1983 value. This exercise shows that, first, in the absence of a decline in unions, employees at unionized establishments would have contributed \$6.66 billion, extending unions’ potential influence to an additional \$2.35 billion. Second, the larger number of employees exposed to unions would have induced a larger shift in funds amounting to \$744 million. Thus, the decline in the union membership rate has cost Democratic candidates (relative to Republicans) an amount more than \$232.7 million in donations from employees.

### v5

@@ 6.2 Ideological versus Strategic Motives
As discussed in Section 2.2, campaign contributions can be driven by either expressive (ideological) or investment-oriented (strategic) motives. While a large strand of work on money in politics finds that individual donors allocate their contributions primarily based on ideological preferences (Stone2010; Bonica2014; Bonica2018; Barber2016; Barber2016a), more recent research has found that parts of employees’ donations are also related to their employers’ strategic interests (Stuckatz2022; Teso2023). Along that line, the observed shift in managers’ donations from Republican to Democratic candidates could reflect, for example, a strategic signaling of goodwill and understanding of unions’ and workers’ interests instead of a true shift in managers’ preferences. Moreover, unionization could change workers’ efforts to support candidates in exchange for future legislative votes on worker-friendly policies. While it is difficult to fully disentangle donors’ motives, in the following we present several pieces of evidence that all point toward changes in ideological preferences being the main driver of our results.
@@
Differentiating candidates by within-party ideology. To examine the ideological patterns in campaign contributions in more detail, we study the changing support for ideologically different candidates within the same party. Kuziemko2023 document the growing influence of conservative factions within the Democratic Party that place less emphasis on predistributive economic policies, such as higher minimum wages, protectionism, and strong labor unions. They trace the origins of this shift to the 1970s, coinciding with the period in which unionization rates began to decline more sharply. To analyze how unionization affects the support of different party factions, we make use of Bonica2014’s (2014) CF scores that assign each recipient an ideal point along a liberal-conservative scale. Democratic candidates are defined “liberal” (as compared to “moderate”) if their CF score lies below the median CF of all Democrats observed in our sample of matched contributions. Similarly, we distinguish between “conservative” and “moderate” Republicans using the median Republican CF score. Table 4 shows the DiD results for the effect on the amount contributed to each of the candidate types. Considering first all employees jointly, we see strong differences in the effects of unionization by the within-party ideological positions of candidates. Unionization significantly increases employees’ support for the most liberal Democrats and decreases support for the most conservative Republicans. In contrast, contributions to moderate Democrats or Republicans are not significantly affected. These results are similar when we focus on donations from managers only, and also for workers the increased support for Democrats is more pronounced for more liberal Democrats. Overall, our results do not seem to be driven by a mere signal of affiliating with the Democratic Party but instead reflect a shift in contributions between clearly distinguishable conservative and liberal candidates.
@@
Differentiating candidates by union support. We continue by examining whether the partisan shift in employee contributions can be solely explained by a shift from candidates not supported to candidates supported by the union organization. To identify union-supported candidates, we match donations from PACs associated with the union organization that was on the ballot in a given union election (including PACs of local union branches). Appendix Table A.11 reports for each union organization in our sample the share of matched contributions to Democratic (as opposed to Republican) candidates. On average, union PACs give 94% of their donations to Democrats, which demonstrates the strong alliance between labor unions and the Democratic Party (Dark2001). In Table 5, we use this information to estimate the effects of unionization on employees’ contribution amounts to candidates supported and not supported by the respective union in a given cycle. Unions may encourage their members to contribute to the same candidates as their PAC, either directly through providing information about these candidates or indirectly through employees contributing to the union PAC and observing which candidates are supported by the union PAC. While the estimated coefficients for the effect of unionization on employees’ contribution amounts to union-supported candidates are positive, they are small and insignificant. [^41] When taking the difference in the amounts going to union-supported and non-union-supported candidates, the effect of unionization is significant only for contributions from managers and only at the 10% level (not shown). Moreover, we find a shift in contributions from Republican to Democratic candidates that remains significant even after conditioning on whether or not candidates are supported by the union. In other words, among both union-supported and non-union-supported candidates we observe significant effects on contribution amounts to Democratic relative to Republican candidates. These results suggest a limited role for unions’ direct influence in channeling employees’ contributions toward candidates in which the union has strategic interests. Instead, unionization appears to cause a more general leftward shift in employees’ donation behavior.
@@
Contributions to PACs. In Table 2, we have shown that contributions to PACs account for a large share of employees’ political contributions. Besides changing support for specific candidates, unionization may also affect employees’ donations to intermediary committees with varying ideological positions. Most evidently, unions may solicit contributions from their members to their own PACs. Table 6 reports DiD estimates for the effect of unionization on employees’ donations to party PACs and interest group PACs, where the latter are further disaggregated into labor union, membership organization, corporate, and trade association PACs. Besides considering the total amount given to these committees, we also measure partisan support by the difference in contribution amounts to Democratic versus Republican PACs. For interest group PACs, party affiliation is determined from the recipients of the PAC’s own campaign contributions. [^42] To track contributions that PACs donate themselves, we exploit that Bonica2019 has matched recipient identifiers to contributor identifiers for recipients’ own contributions. Based on the matched outgoing contributions from PACs, we define an interest group PAC as “Democratic” (“Republican”) if more (less) than 50% of its campaign contributions goes to Democratic candidates in a given election cycle. Considering first the contributions from all employees of an establishment to party PACs, the results mimic those for candidate contributions. While there is no effect on total amounts, unionization leads to a significant shift from Republican to Democratic committees. Among interest group PACs, there is a significantly positive effect on donations to union PACs and a significantly negative effect on donations to corporate PACs. When distinguishing between donations from workers and managers, results differ somewhat. For workers, we see a significant increase in the total amounts donated to both party and interest group PACs, which implies that unions are successful in mobilizing PAC contributions from workers. The increase in donations appears to be driven by unions and membership organizations, pointing toward increased support for labor and civil society interest groups. In contrast to our results on candidates, however, we do not see a significant shift across party affiliations. For managers, the results are similar to those on candidate contributions. While there is no effect on overall party PAC spending, managers increasingly donate to Democratic rather than Republican PACs. Moreover, donations to corporate PACs drop, which highlights that unionization can decrease managers’ support for business interest groups. [^43] Note that the absence of significant effects on managers’ donations to union PACs, in contrast to the significant increase observed among workers, provides supporting evidence for our occupational classification. If individuals classified as managers were in fact misclassified union-eligible workers, we would expect to observe increased donations to union PACs among them as well. The lack of such an effect suggests that our manager group largely consists of individuals who are correctly identified as excluded from unionization. Overall, these results match with the observed pro-liberal shift in workers and managers’ contributions to political candidates.
@@
Contributions from firm PACs. Finally, we directly examine whether unionization influences strategic donations of unionized firms. A growing body of evidence suggests that corporations allocate their PAC contribution in ways that are consistent with an attempt to buy access to relevant politicians (Kalla2016; Powell2016; Fouirnaies2018). For each union election in our estimation sample, we link firm PAC contributions given to candidates and aggregate them at the firm level. [^44] We link firm PAC contributions using a fuzzy match of the firm name. Given that the location of the firm PAC may differ from that of the establishment, we do not match on the commuting zone of the establishment. For 1,902 out of the 6,603 establishments in our estimation sample, we observe at least one firm PAC contribution over our observation window. In Appendix Table A.12, we report results for the full sample of establishments when considering unmatched observations as zeros, as well as for the reduced sample of establishments with at least one matched firm PAC contribution. The results of our DiD model, reported in Appendix Table A.12, show that unionization does not significantly affect firm PACs’ total contribution amounts or their partisan composition. We conclude that the leftward shift in employees’ donations does not coincide with a corresponding shift in employers’ strategic donations.
6.3 Labor Relations Environment
Our findings indicate that unionization increases the alignment of workers’ and managers’ political preferences, shifting support toward Democratic candidates. This influence can operate through an information channel, where unions provide political messaging that persuades both workers and managers, but it may also stem from shifts in workplace dynamics. If unionization facilitates cooperative contact between workers and managers, it may promote perspective-taking and foster managers’ understanding of workers’ political positions (Allport1954). The contact hypothesis literature suggests that interactions between different groups are most likely to reduce cleavages when they occur on an equal footing, involve cooperation toward common goals, and are governed by mutually accepted norms and regulations (pettigrew2006).
Case studies from health care and manufacturing underscore how these dynamics can unfold in practice. At Kaiser Permanente and in joint ventures like NUMMI (Toyota-GM) and Saturn (GM-UAW), labor was embedded in systems of shared governance, including joint committees, co-developed performance metrics, and continuous problem-solving mechanisms (e.g., BrownReich1989; Rubinstein2000; Kochan2008). These arrangements closely reflect the conditions described by contact theory, fostering ongoing interaction under cooperative norms. Consistent with this, political behavior at Kaiser Permanente shifted soon after the establishment of its labor–management partnership in 1997. Individual political contributions from employees were nearly evenly split in the late 1990s but became overwhelmingly Democratic—reaching 83% by 2004 and nearly 90% thereafter (OpenSecrets2025). At the same time, institutional design alone does not seem to guarantee such outcomes. GM’s Van Nuys plant, which sought to replicate NUMMI-style collaboration, reveals the fragility of such efforts in the absence of trust and mutual commitment. Despite adopting similar team-based structures, the initiative faltered amid adversarial labor relations and management’s reluctance to cede meaningful authority (BrownReich1989). In that setting, formal cooperation did little to foster genuine perspective-taking, suggesting that the political consequences of unionization depend not only on structural features, but on the quality and depth of the interaction itself.
To investigate whether workplace contact can help explain the shifts in campaign donations in response to unionization, we examine how effects vary across different labor relations environments.
Right-to-Work laws. As a first measure of labor relations climate, we study the role of right-to-work (RTW) legislation. RTW laws allow workers in unionized establishments to forgo union membership and dues payments while still benefiting from union representation, thereby encouraging free-riding of workers and reducing unions’ financial resources. RTW measures are typically advocated for by conservative activists as a means of reducing union strength and have been most successfully adopted in Southern U.S. states with more pro-business policies (holmes1998; rao2011) and conflicts between unions and employers (dixon2008; haskett2024). Enacting RTW laws is indeed consequential: at the local level, it is found to reduce union membership rates (ellwood1987; eren2016) and depress voter turnout and vote shares for Democratic candidates (Feigenbaum2018). We complement this evidence by studying how RTW laws moderate the effect of unionization on employees’ campaign contributions at the establishment level.
@@ 6.3 Labor Relations Environment
Our results confirm that the political effects of unionization are significantly stronger in non-RTW states. In Table 7, Panels A.1 and A.2, we split our estimation sample based on whether or not the union election takes place in a state that has an RTW law in force at the time of the election. In states without RTW laws, we see significantly positive effects of unionization on support for Democratic over Republican candidates, while for RTW states the coefficients are smaller and not significant. This is true for all employees, as well as for workers and managers separately. Thus, unions’ ideological influence on employees seems to be substantially lower in states with reduced union power and more adversarial labor relations.
Unfair labor practice charges. A second factor shaping union-employer relations is the extent of employer opposition to unionization. A long-standing literature documents the widespread use of anti-union campaigns by employers in the United States, often involving aggressive tactics to deter unionization (Freeman1990a; Kleiner2001; Logan2002; Schmitt2009). One measure of employer opposition is whether a union election is accompanied by a charge for unfair labor practices (ULPs) against the employer. The NLRA allows workers to file ULP complaints when employers interfere with their rights to organize and collectively bargain (Bronfenbrenner2009; McCammon2001; Stansbury2021). [^45] These interferences can include restraining employees in their rights to organize in a union or collectively bargain (NLRA § 8(a)(1)), dominating or controlling a union (§ 8(a)(2)), discharging or otherwise discriminating workers involved in organizing (§§ 8(a)(3) and 8(a)(4)), and failing to bargain in good faith with the union (§ 8(a)(5)). To focus on employer misconduct in the process of the organizing drive, we restrict our analysis to charges for violations of §§ 8(a)(1), 8(a)(3), and 8(a)(4). We gather and merge data on ULP charges against employers that were filed to the NLRB between 1984 and 2020.464646The data stems from the CHIPS archive for 1984-2000, from the CATS archive for 1999-2011, and from the NLRB website for 2007-2020. Since charges must be filed within six months of an alleged violation, we consider all ULP charges that were filed between six months before to six months after a union election. See Appendix B.4 for more details on the data preparation. In our estimation sample, 44% of all union elections involve a ULP charge.
Our results in Panels B.1 and B.2 of Table 7 reveal that the political effects of unionization are substantially weaker in establishments where ULP charges are filed. In union elections without ULP charges, we observe a large and statistically significant increase in contributions to Democratic over Republican candidates among both workers and managers. However, when elections involve ULP charges, the estimated effects are smaller and remain marginally significant only for workers. These findings suggest that when unions must combat aggressive employer opposition, they may struggle to exert political influence on employees. When managers strongly resist unionization, the potential for cooperative contact and perspective-taking may diminish, reducing the likelihood that managers shift their political preferences in response to unionization.
Collective bargaining contracts. Finally, we study heterogeneity by whether the union and employer reach an agreement after a union election is won, in particular whether they successfully negotiate a collective bargaining contract. Collective bargaining is a defining feature of unionized workplaces, yet not all successful unionization drives are characterized by cooperative labor-management interactions that lead to an agreement. Compiling and merging data on contract negotiation notices from the Federal Mediation and Conciliation Service (FMCS), we only observe a bargaining contract for 48% of the won union elections. [^47] We combine contract data compiled by Holmes2006 for 1985-2003 and by Gregg2024 for 2004-2020. The data includes notices of contract expirations that are provided to the FMCS for it to be able to prepare mediation services for the renegotiation of a contract. We consider contract expiration notices in the five years following the union election. See Appendix B.4 for more details on the data preparation.
We find strong political effects of unionization only in establishments where a collective bargaining agreement is reached. In Panels C.1 and C.2 of Table 7, we divide our treatment group into won union elections with a reported bargaining contract and those without. [^48] In both cases, we compare the won union elections to the control group of all lost union elections. For union certifications that result in a collective bargaining agreement, both workers and managers exhibit a significant shift toward supporting Democratic over Republican candidates, while we find no significant effects in establishments without a bargaining agreement. For workers, the results may suggest that organizing alone is not enough – unions must also deliver on the bread and butter issues to be able to politically influence their members. For managers, the results are particularly noteworthy. Although they do not directly benefit from a collective bargaining contract, they nonetheless exhibit a larger ideological shift in their donations in establishments where an agreement is reached. This result is again consistent with the idea that sustained, cooperative contact – during successful contract negotiations – can help explain why managers align their political positions with workers and the union.